A pressure of a gas container varies directly with its temperature and inversely with its volume. At a temperature of 70 Kelvin and a volume of 15 cubic meters, the pressure is 19.6 newtons per square meter. What is the pressure if the temperature is increased to 80 Kelvin and the volume is decreased to 12 cubic meters?
Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

A pressure of a gas container varies directly with its temperature and inversely with its volume.
p = k*t/v
------
Solve for "k" using "At a temperature of 70 Kelvin and a volume of 15 cubic meters, the pressure is 19.6 newtons per square meter."
19.6 = k*70/15
70k = 294
k = 4.2
-----
Equation:
p = 4.2*t/v
What is the pressure if the temperature is increased to 80 Kelvin and the volume is decreased to 12 cubic meters?
p = 4.2*80/12
p = 21 newtons per sq meter
------------------------------
cheers,
Stan H.
---------------

Answer by richwmiller(17219)

p=k*t/v
19.6=k*70/15
15*19.6=k*70
(15*19.6)/70=k
k=4.2
p=4.2*80/12
p=4.2*20/3
p=1.4*20
p=28